OpenStudy (anonymous):

Find le slope and y-intercept of 3x-5=0

OpenStudy (anonymous):

What the freckles do I do? There's no y and I'm so confuzzled by this... :P

OpenStudy (anonymous):

Cuz' normally the formula is y=mx+b

OpenStudy (anonymous):

m=______ b=______

OpenStudy (anonymous):

make y equal the expression that is set equal to 0

OpenStudy (anonymous):

...what?

OpenStudy (anonymous):

okay, so you can graph the equation as y = 3x -5

OpenStudy (anonymous):

But isn't there a zero?

OpenStudy (anonymous):

m=3 and b=-5 now...

OpenStudy (anonymous):

y = 3x - 5 is the same as y + 0 = 3x - 5 + 0

OpenStudy (anonymous):

but what if the equation is x=2? I don't understand it.

OpenStudy (anonymous):

Does it become y=-x+2

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

m=-1 b=2?

OpenStudy (anonymous):

uh huh

OpenStudy (anonymous):

so the other one...

OpenStudy (anonymous):

Ohkay, thankyouu :3

OpenStudy (anonymous):

you're welcome